A titration of pure vitamin C was performed by titrating a sample of vitamin C solution with 23.35 mL of a 0.43 M Iodine solution. If there was 20.0 mL of vitamin C in the flask, what was the initial concentration of the vitamin C solution? The vitamin C solution was not diluted in anyway.
1
Expert's answer
2019-09-27T04:46:46-0400

From titration equation-

"M_1V_1=M_2V_2"

Where

"M_1= 0.43\ M \ iodinesolution\\V_1=23.35 \ ml\\V_2=20\ ml"

"M_2=\frac{M_1V_1}{V_2}=\frac{0.43\times 23.35}{20}=0.50 \ molar"

So initial concentration of vitamin C was "0.5\ M"